40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 182     Entry time: Thu Apr 17 21:39:25 2014
Author: Koji 
Type: Optics 
Category: General 
Subject: More alignment 

STORY:

- The cavity mirrors have scattering spots. The cavity alignment should have been scanned to find a cavity mode to have lowest loss possible.
  BTW, We only have horizontal dof for the alignment scan.

- After some struggle nice cavity mode was found. The cavity transmission was 96% for the ideally matched TEM00 carrier.

- It turned out that this imposed too much beam shift in the input beam (~2mm).

- This big shift induces a lot of trouble for the peripheral optics (PDs, QPDs, sterring mirrors).

- What should we do???

Analysis:

- The beam needed to go up between CM1 and CM2 to have the right spots on them. ("UP" is the input side of the OMC).

- This imposed the beam between FM1 and FM2 moved up. In other word, for the given alignment of the FMs by the template,
  We needed to hit the upper part of the FMs to have the spots on the CMs up.

Solution:

- The above argument suggets that the nominal beam will give us the right spots on the CMs if we rotate the FMs.
  Of course this induces the spot move on the FMs. But this should not be the issue as the most of the loss seems to come from the CMs.

- How much misalignment show we give to the FMs? We want to shift the beam by 2mm on the CMs.
  The length of the optical lever is ~0.25m. Therefore the mialignment angle should be

  theta = 2e-3/2/0.25 = 4e-3 rad = 4mrad.

  The template pad has ~20mm separation. The thickness of the shim should be 20mm*4mrad = 80um

- Our aluminum foil seems to have the thickness of 30-40um. We can't have this minimum thickness on the template pad as there is not enough compression pressure
  => Just use a single layer of Al piece to shim the FMs.

Attempt:

- The shims were inserted at the upper pads of the FMs.

- Aligned the input beam and the CMs so that the spots on the CMs are approximately recovered.

- Measure the cavity power budget

Pin: 34.7mW
Refl PD: offset = -7.5mV, unlock = 6.07V, inlock = 89.7mV

Ptrans = 32.5mW

Ptrans(CM2) = 0.181mW
Ptrans(CM2) = 0.184mW

Assume finesse of 400

==>

Pin: 34.7mW
Pjunk: 0.534mW
Pcoupled: 34.1mW

Mode matching: 98.5%

Cavity reflectivity in power: 0.00061
Cavity transmission in power: 0.951 (This is not a best number but acceptable.)
Loss per mirror: 75.4ppm

FM power refl/trans: 0.9923 / 7630ppm
CM1 power refl/trans: 0.999882 / 42.8ppm
CM2 power refl/trans: 0.999881 / 43.5ppm
Total roundtrip loss of the cavity (Loss + CM leakage): 388ppm

Result:

How much the input beam is away from the left wall of the OMC breadboard?

40.88mm from the template edge
  8.36mm between the template edge and the bread board
=> 32.52mm

How much should this number be? 32.94mm from the solidworks model => With in 0.5mm! Nice!

Next:

- Just in case plce all of the optics and check if the beam is delivered within the alignment range of the optics

 

ELOG V3.1.3-